### v1

# Teager-Kaiser Energy Methods for EEG Feature Extraction in Biomedical Applications

Abstract
Electroencephalography (EEG) signals are inherently non-linear, non-stationary, and vulnerable to noise sources, making the extraction of discriminative features a long-standing challenge. In this work, we investigate the non-linear Teager–Kaiser Energy Operator (TKEO) for modeling the underlying energy dynamics of EEG in three representative tasks: motor imagery, emotion recognition, and epilepsy detection. To accommodate the narrowband nature of the operator, we employ Gabor filterbanks to isolate canonical frequency bands, followed by the Energy Separation Algorithm to decompose the TKEO output into amplitude envelope and instantaneous frequency components. We then derive a set of energy descriptors based on this demodulation and compare their classification performance against established signal energy and power spectrum features. TKEO features outperform the respective baselines in motor imagery and epilepsy detection, whereas they perform on par in emotion recognition. Our findings suggest that the proposed TKEO-based pipeline provides an intuitive framework for extracting EEG signal dynamics.
footnotetext: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
Index Terms—  Teager-Kaiser Energy Operator, Amplitude–Frequency Modulation, EEG, Biomedical Applications
1 Introduction
Human brain activity underlies critical cognitive and affective processes, yet many aspects of its neural mechanisms are not adequately understood. Among the available recording choices, Electroencephalography (EEG) is particularly popular, as it offers cost-effective, minimally invasive means of capturing oscillatory neural dynamics at high temporal resolution. At the same time, EEG signals are high-dimensional and notoriously noisy, which makes the extraction of informative features a persistent challenge. Despite the success of deep learning approaches in other domains, analysis of biosignals like EEG is ultimately bounded by limitations in data availability and subject heterogeneity, which has led many studies to continue relying on handcrafted, domain-specific features rather than purely data-driven representations [malekzadeh2021]. Conventional spectral and energy-based measures mainly capture overt oscillatory power [psd1], often overlooking the transient and non-linear dynamics that govern several neurophysiological behaviors.
Traditional EEG analysis has relied on handcrafted features designed to capture statistical, spectral, and dynamical signal properties. Statistical descriptors such as mean, variance, zero-crossing counts, and autocorrelation have long been used for baseline characterization [EEG-featutres_stats, EEG-featutres_zc, EEG-featutres_autocorrelation, EEG-featutres_rms_autocorrelation]. However, spectral approaches, particularly Power Spectral Density (PSD), remain the most common approach for quantifying canonical oscillatory rhythms [psd1, psd2]. Recently proposed complementary measures to quantify temporal and spatial complexity include differential entropy [invastigating_critical], and asymmetry indices [EEG-featutres_asymmetry], respectively. Multiscale and non-stationary dynamics are effectively captured by Discrete Wavelet Transform (DWT) and Empirical Mode Decomposition (EMD) [emd_dwt, dwt, emd], while non-linear descriptors, including fractal dimensions [avramidis2021] and higher-order spectra [higher_order_spectra] have been used to capture the chaotic structure of EEG. Despite this rich toolbox, most approaches rarely target an explicit analysis of transient and non-linear dynamics in EEG.
To address this limitation, the Teager–Kaiser Energy Operator (TKEO) [kaiser, maragos1993] provides a complementary approach. As a non-linear differential operator, TKEO yields instantaneous energy by estimating modulations of both amplitude and dominant frequencies in oscillatory signals. This makes it well-suited for EEG analysis, where recorded oscillatory activity can be naturally decomposed into frequency bands with distinct clinical functions [eeg_basics_book].
TKEO has been employed across signal processing domains for detecting instantaneous energy fluctuations in oscillatory signals. In speech processing, it has been leveraged for tasks such as formant detection [tkeo_formants], denoising [speech_denoising], replayed speech identification [replay_attacks], and emotion recognition [speech_emotion, speech_emotion2]. In music signal analysis, TKEO has been used for tempo estimation [music_beat_tracking], note onset detection [music_note_detection], instrument  [zlatintsi2012] and genre recognition [zlatintsi2014]. Beyond acoustic signals, TKEO has proven effective in biomedical applications to extract physiologically plausible energy fluctuations. In cardiovascular research, for example, it has enabled automated phonocardiogram (PCG) analysis for the detection of heart valve disorders without requiring prior segmentation [detection-of-heart-disorder]. With respect to analysis of neurophysiological recordings, TKEO has been used for electromyography (EMG) segmentation to detect precise movement onsets [upper_limb_movement], while modified thresholding strategies have further improved detection accuracy [stkeo]. TKEO has also been applied to EEG, including differentiating sleep states [sleeping_periods], detecting focal versus non-focal activity [focal], as well as epileptic activity [epilepsy_detection].
In this study, we systematically evaluate TKEO-based features across two tasks characterized by well-established transient dynamics—motor imagery and epilepsy detection—as well as in emotion recognition, which tests the method’s applicability beyond its original motivating context. Through this evaluation, we clarify the strengths and limitations of TKEO features in EEG analysis and identify conditions under which it provides added value. While end-to-end deep learning is increasingly applied to EEG analysis, our methodology emphasizes model comparisons and feature interpretability, elements that are obscured and difficult to quantify in end-to-end training. TKEO-based energy descriptors instead improve upon baseline performance in motor imagery and epilepsy detection while still performing on par in emotion recognition. Our contributions can be summarized as follows:
  • We motivate and extract a comprehensive set of EEG features derived from instantaneous energy analysis using TKEO.
  • We introduce a systematic evaluation of TKEO-derived features against conventional PSD-based energy and squared-energy features across three representative EEG benchmarks.
  • We provide empirical evidence of the conditions under which TKEO features outperform and/or yield complementary insights over alternative energy measures, particularly in tasks including transient and non-linear dynamics.
2 Methodology
2.1 Teager-Kaiser Energy Operator
Many physiological signals exhibit oscillatory electrical activity that can be described in terms of amplitude and frequency modulations (AM-FM). To capture such modulations, TKEO was introduced as a non-linear differential operator that estimates the instantaneous energy of a signal [kaiser, maragos1993]. This operator estimates the instantaneous energy of a signal as a function of its time-varying amplitude and frequency, effectively serving as a tracker of energy variations. TKEO was originally formulated for both continuous and discrete signals. For a continuous signal $x(t)$ it is defined as
$\Psi_{c}[x(t)]=(\dot{x}(t))^{2}-x(t)\,\ddot{x}(t),$ (1)
while for a discrete-time signal $x[n]$ it takes the form
$\Psi_{d}[x[n]]=x^{2}[n]-x[n-1]\,x[n+1].$ (2)
For the purpose of this study, we constrain our analysis to the discrete-time TKEO formulation, i.e., $\Psi\equiv\Psi_{d}$.
2.2 Filterbank Decomposition
TKEO provides reliable instantaneous energy estimates primarily when applied to narrowband signals [maragos1993]. To apply this premise, the use of filterbanks helps decompose the broadband (raw) signal into localized frequency subbands, which, as a result, produces smoother instantaneous amplitude and frequency modulations [dimitriadis2004]. Among possible choices, Gabor filters are particularly suitable due to their joint time–frequency resolution and compactness [maragos1993].
A Gabor filterbank is constructed by generating a sequence of Gabor filters whose center frequencies are uniformly distributed between a lower ($f_{\mathrm{low}}$) and an upper ($f_{\mathrm{high}}$) cutoff frequency. In this way, the broadband signal is decomposed into subbands, each filter capturing localized spectral content around its center frequency. In the EEG setting, we apply this procedure to both the broadband signal (0.5-100 Hz) and each canonical frequency band: Delta (0.5–3 Hz), Theta (4–7 Hz), Alpha (8–12 Hz), Beta (13–30 Hz), and Gamma (30–50 Hz). For each band, $f_{\mathrm{low}}$ and $f_{\mathrm{high}}$ are set to the band boundaries, and $N$ filters are uniformly distributed inside this interval. The frequency step is defined as: $\Delta f=\frac{f_{\mathrm{high}}-f_{\mathrm{low}}}{N}$, where $N$ is the number of filters. The center frequencies are then: $f_{\mathrm{c}}(k)=f_{\mathrm{low}}+k\Delta f$, where $k\in\mathbb{Z}_{\geq 0}:k<N$. Once the set of $f_{c}$ has been determined, each EEG signal is filtered with the corresponding Gabor filters, and TKEO is applied to each filter’s output. This results in $N$ energy signals, one for each subband.
For each TKEO signal, we compute the temporal mean as a measure of the average subband energy. Finally, for both the broadband and each canonical band, the filter of maximum averaged TKEO value is selected, and the corresponding narrowband signal is retained for further analysis, as it represents the subband in which the TKEO provides the most informative characterization.
2.3 Energy Separation Algorithm
To estimate the amplitude envelope and instantaneous frequency of AM–FM signals, Maragos et al. [maragos1993] introduced the Energy Separation Algorithm (ESA), which employs non-linear combinations of TKEO outputs to decouple amplitude and frequency components. For discrete-time signal $x[n]$, this framework was extended to the Discrete Energy Separation Algorithm (DESA) [maragos1993]. Here we adopt the DESA-1 variant, for which the amplitude envelope $|a[n]|$ and the instantaneous frequency $\Omega[n]$ of $x[n]$ are defined as
$|a[n]|\approx\sqrt{\frac{\Psi[x[n]]}{1-\left[1-\frac{\Psi\big[y[n]+y[n+1]\big]}{4\,\Psi[x[n]]}\right]^{2}}},$ (3)
$\Omega[n]\approx\arccos\!\left(1-\frac{\Psi\big[y[n]+y[n+1]\big]}{4\,\Psi[x[n]]}\right),$ (4)
where y[n]=x[n]x[n1]y[n]=x[n]-x[n-1]. An example of estimating the amplitude envelope and the instantaneous frequency, from an EEG signal is depicted in Fig. 1. These quantities are subsequently used for feature extraction, as described in Sec. 2.4.
2.4 TKEO Feature Extraction
We extract TKEO-based features from each EEG channel across either the predefined frequency bands, or the broadband signal. We first compute the mean Teager–Kaiser energy (m-TKEO). For a band-limited signal $s$, the relative contribution of its Teager–Kaiser energy with respect to all frequency bands under consideration is:
$\text{RE}_{\,\text{band}}(s)=\frac{\Psi_{\text{band}}(s)}{\sum_{k}\Psi_{k}(s)},$ (5)
where band $k\in\{\text{Delta},\text{Theta},\text{Alpha},\text{Beta},\text{Gamma}\}$. Consequently, we calculate the mean Relative Energy (m-RE). Finally, applying DESA-1, we obtain the instantaneous amplitude $|a(n)|$ and frequency $\Omega(n)$, from which we compute the mean Instantaneous Amplitude Modulation (m-IAM) [zlatintsi2012] and the variance of Instantaneous Frequency Modulation (v-IFM). From this point onward, by TKEO features we refer to the following 4D feature set: m-TKEO, m-RE, m-IAM, and v-IFM features.
(a) Amplitude Envelope
(b) Instantaneous Frequency
Fig. 1: Estimated demodulated signal of the C3 EEG channel from a sample recording of the BCI-IV 2a dataset.
3 Experimental Setup
3.1 Datasets
BCI Competition IV Dataset 2a: The BCI-IV 2a dataset [bci2a] consists of EEG recordings from 9 subjects performing 4 motor imagery tasks, i.e., left-hand, right-hand, feet, and tongue movement imagery. Each subject completed two sessions on different days, with each session comprising 6 runs of 48 trials (12 per class), resulting in 288 trials per session. Each trial lasted 6 seconds in total, with an onset cue presented after the first 2 seconds. Subjects performed the instructed motor imagery for the subsequent 4 seconds, which was used as the analysis window. EEG was acquired from 22 channels
Table 1: Dataset-wise preprocessing parameters
Dataset Window (s) Overlap (%) # Filterbanks
BCI-IV 2a 4 0 25
SEED 20 50 25
TUEP 10 0 12
following the 10–20 configuration [10-20system], along with 3 electrooculogram (EOG) channels, all sampled at 250 Hz.
SEED: The SEED Dataset [invastigating_critical, seed2] is a widely used benchmark for EEG-based emotion recognition. It comprises EEG recordings from 15 participants while they viewed 15 Chinese film clips, each lasting approximately 4 minutes. Each video was selected to induce one of three affective states –positive, neutral, or negative– resulting in a single categorical label per clip. The EEG signals were recorded using a 62-channel setup following the 10–20 configuration, with an original rate of 1000 Hz, subsequently downsampled to 200 Hz.
TUH Epilepsy Corpus: The TUH Epilepsy Corpus (TUEP) [tuh] is a curated collection of clinical EEG data from Temple University Hospital, including recordings from matched 100 epileptic and 100 healthy subjects. In total, TUEP contains 698 sessions, yielding 1,785 EEG recordings from epileptic and 513 from non-epileptic patients. Individual recording durations vary from a few seconds to approximately one hour. All EEG signals were acquired using the 10–20 configuration with 19 channels and sampled at 250 Hz with supplementary electrocardiogram (EKG) channels also available.
3.2 Preprocessing
All recordings were segmented into overlapping windows, where window length and overlap ratio were determined from the characteristics of each individual dataset, as summarized in Table 1. Out of the three datasets, SEED was provided in a preprocessed format, which included manual removal of EMG and EOG artifacts, notch filtering at 50 Hz and lowpass filtering at 75 Hz [invastigating_critical]. For the rest of the datasets, we applied a similar pipeline that involved Independent Component Analysis (ICA) to remove EKG artifacts in TUEP and EOG artifacts in BCI-IV 2a. Then, notch filtering was applied at 60 Hz for TUEP and 50 Hz for BCI-IV 2a to eliminate powerline interference. Finally, all signals were highpass-filtered at 0.5 Hz.
Before TKEO-based feature extraction, each EEG segment was decomposed into five narrowband components using bandpass filtering corresponding to the canonical EEG bands. Within the frequency
Table 2: 5-fold Mean Classification Performance and Standard Deviation on BCI-IV 2a.
Features Balanced Accuracy (%) ROC-AUC ($\mathbf{\times 10^{2}}$)
Raw Fused Raw Fused
[…]
range of each canonical band we applied filterbanks of 25 filters (SEED, BCI-IV 2a) and 12 filters (TUEP, due to computational constraints), see also Table 1.
3.3 Baseline Feature Extraction
To enable a comparative evaluation of the proposed approach and to assess its potential complementarity with established EEG descriptors, we re-implement a standard feature extraction pipeline for EEG signals, wherein we first isolate the desired frequency bands with a 10th-order Butterworth filter, and then extract a set of baseline features from the isolated bands (or the raw signal). These features include the Power Spectral Density (PSD) and the Signal Energy of the signal. PSD is computed using the Welch method, yielding $f_{s}/2$ features per signal, where $f_{s}$ denotes the sampling frequency. The Signal Energy, defined as the squared magnitude of the raw signal at any point in time and commonly used in signal processing as a direct measure of instantaneous energy, is employed here. Similar to the proposed TKEO statistical descriptors, the mean Signal Energy and the mean Relative Signal Energy features are extracted.
For clarity, we define the term Combined features as the set consisting of the TKEO features, the PSD, mean Signal Energy, and mean Relative Signal Energy, concatenated.
3.4 Evaluation Protocol
We evaluate the utility of TKEO-based features on three classification tasks: motor imagery (BCI-IV 2a dataset), emotion recognition (SEED dataset) and epilepsy detection (TUEP dataset). A Subject Independent setting is presented, where models are trained on a subset of subjects and tested on unseen individuals. This setting reduces the risk of confounding due to subject- or recording-specific biases and better assesses cross-subject generalization. While we also conducted experiments in a Subject Dependent setting (i.e., training and testing on different trials from the same subject), we report only the independent results here, as they better reflect real-world applications and due to space constraints.
Features from all EEG channels were concatenated and the resulting feature vectors were standardized with a Standard Scaler and classified using a Random Forest classifier with 100 estimators, chosen for computational efficiency. We deliberately employed Random Forest, as it is a standard classifier that allows us to assess the discriminative power of the extracted features. This choice ensures that classification performance reflects the quality of the proposed TKEO-based features, rather than the capacity of complex machine learning models. Evaluation is performed with stratified 5-fold cross-validation. SEED and BCI-IV 2a are balanced, whereas TUEP is imbalanced, with 77.7% epileptic recordings; to mitigate this, we report metrics robust to class imbalance, namely Balanced Accuracy and ROC-AUC.
4 Results & Discussion
Table 3: 5-fold Mean Classification Performance and Standard Deviation on SEED.
Features Balanced Accuracy (%) ROC-AUC ($\mathbf{\times 10^{2}}$)
Raw Fused Raw Fused
[…]
Table 4: 5-fold Mean Classification Performance and Standard Deviation on TUEP.
Features Balanced Accuracy (%) ROC-AUC ($\mathbf{\times 10^{2}}$)
Raw Fused Raw Fused
[…]
4.1 Comparison to Baselines
The classification results for all features are summarized in Tables 24. We report results using features extracted directly from the raw signal (raw-band) as well as from the fusion of features computed across all canonical frequency bands (fused-band). TKEO features generally outperform the baseline features (PSD and Signal Energy) in motor imagery and epilepsy detection tasks, whereas in emotion recognition, the baselines achieve higher performance.
As shown in Table 2, TKEO features consistently outperform all other feature sets in the motor imagery task, with the fused-band achieving the best overall performance (39.8% Balanced Accuracy and 66.2% ROC-AUC), substantially exceeding chance level. Moreover, integrating TKEO features with baselines yields performance gains over the baselines alone in most cases, highlighting their complementary nature. In the SEED dataset (Table 3), baselines generally outperform TKEO features, except for the raw-band ROC-AUC (82.0%) of Combined set, reflecting strong class separability. In the TUEP dataset (Table 4), Combined features consistently achieve the highest performance. In particular, the fused-band configuration achieves the highest Balanced Accuracy (76.2%)—approximately 15% higher than other feature sets—and the highest ROC-AUC (82.2%). TKEO features alone attain performance comparable to the baselines, further highlighting their complementary contribution when combined.
Overall, TKEO features show stronger relative performance in motor imagery and epilepsy detection compared to emotion recognition, underscoring their suitability for tasks where transient temporal changes are important and their limitations when the discriminative information is not as temporally localized. Additionally, the fused-band consistently outperforms the raw-band within TKEO features, reflecting the advantage of aggregating information across frequency bands while preserving the narrowband characteristics of TKEO.
4.2 Ablation Study
In our study, EEG signals were decomposed into frequency bands using a filterbank with a fixed number of filters. To examine the impact of the number of filters, we performed an ablation study on the BCI-IV 2a dataset, repeating the analysis with varying filter counts. The results in Table 5 show that the 50-filter configuration achieves the overall highest performance, yielding a Balanced Accuracy of 41.0% and an ROC-AUC of 66.8% in fused-band. This suggests that a finer filterbank provides higher spectral resolution, enabling the TKEO to capture narrowband EEG dynamics that are otherwise averaged out with fewer filters. To further expand our study, we investigate the impact of each frequency band to the fused-band overall accuracies. As shown in Fig. 2, the SHAP [shap] heatmap highlights elevated contributions in sensorimotor regions (C3, C4, CP3, CP4) in Alpha–Beta bands that capture event-related desynchronization/synchronization (ERD/ERS) during imagined movements. This pattern aligns with well-established neurophysiological findings of motor imagery [scherer2018].
Table 5: 5-fold Mean Classification Performance and Standard Deviation on BCI-IV 2a of TKEO Feature for different Filterbanks.
# Filters Balanced Accuracy (%) ROC-AUC ($\mathbf{\times 10^{2}}$)
Raw Fused Raw Fused
[…]
Fig. 2: Absolute SHAP values of frequency bands indicating their contributions to fused-band TKEO predictions across EEG channels in BCI-IV 2a; higher SHAP values represent greater contribution.
5 Conclusion
In this paper, we investigated the utility of the Teager-Kaiser Energy Operator on three EEG classification tasks, namely motor imagery, emotion recognition, and epilepsy detection. A Gabor filterbank was applied within each canonical band to extract band-specific signals, followed by the application of TKEO to extract statistical features, and instantaneous modulation features using the DESA-1 demodulation algorithm. Experimental results demonstrate that TKEO-based features achieve competitive performance compared to baseline measures, improving accuracy in motor imagery and epilepsy detection tasks, while performing on par with energy-based features on emotion recognition. Feature fusion across frequency bands further improves performance, highlighting the narrowband sensitivity of TKEO. Overall, our analysis indicates that TKEO is a valuable tool for EEG signal analysis, both as a standalone method and as a complementary feature extraction technique. A promising direction for future work is the integration of TKEO into deep learning (DL) frameworks, combining the predictive power of DL with physiologically meaningful, interpretable representations.